Select the option that continues the pattern in this question
3,-6,12,4,20
And
7,16,8,27,9

Found 2 solutions by Edwin McCravy, AnlytcPhil:
Answer by Edwin McCravy(20059)

3,-6,12,4,20

Start with 3

Subtract 9, get -6
Add 18,  get 12
Subtract 8, get 4
Add 16, get 20
Subtract 7, get 13
Add 14, get 27
Subtract 6, get 21
Add 12, get 33
Subtract 5, get 28
Add 10, get 38
Subtract 4, get 34
Add 8, get 42
   
--------------------------

7,16,8,27,9


 7,16,8,27,9,

The odd terms go

 7,  ,8,  ,9,  10,  11,  12,

The even terms go

4*4 = 16
3*3*3 = 27
2*2*2*2 = 16
1*1*1*1*1 = 1
0*0*0*0*0*0 = 0
-1*-1*-1*-1*-1*-1*-1 = -1

Edwin


Answer by AnlytcPhil(1806)

3,-6,12,4,20

Start with 3

Subtract 9, get -6
Add 18,  get 12
Subtract 8, get 4
Add 16, get 20
Subtract 7, get 13
Add 14, get 27
Subtract 6, get 21
Add 12, get 33
Subtract 5, get 28
Add 10, get 38
Subtract 4, get 34
Add 8, get 42
   
--------------------------

7,16,8,27,9


 7,16,8,27,9,

The odd terms go

 7,  ,8,  ,9,  10,  11,  12,  13,...

The even terms go

4*4 = 16
3*3*3 = 27
2*2*2*2 = 16
1*1*1*1*1 = 1
0*0*0*0*0*0 = 0
-1*-1*-1*-1*-1*-1*-1 = -1

7,16,8,27,9,16,10,1,11,0,12,-1,13,...

Edwin